For a saturated solution of CaF2(s) find (F-). Ksp for CaF2(s) = 5.3 x 10-9.

Please show work. I am studying for a test.
Answer #1

Solution:

The solubility product of sparingly salts is calculate from solubility as:

A) For uni-univalent salt: For example AgCl

Ksp = [Ag+] x [Cl-]

B) For uni-bivalent salt: For example. CaF2

Ksp = [Ca2+] [2F-]^2

The dissociation of CaF2 take place as:

CaF2 = Ca2+ + 2F-

S------------S----------2S. (S.= solubilty)

Thus, Ksp = S x (2S)^2 = 4S^3

4S^3 = 5.3 x 10^-9

S^3 = 5.3 x 10^-9 / 4 = 1.325 x 10^-9

S = 1.098 x 10^-3 = 1.1 x 10^-3

Hence, solubility of F- = 2S = 2 x 1.1 x 10^-3 = 2.2 x 10^-3